OpenStudy (anonymous):

I can't get this.. Help? Find the angle between the given vectors to the nearest tenth of a degree. u = <8, 7>, v = <9, 7>

OpenStudy (anonymous):

\[ if~ \theta~be~the~\angle~\between~the~vectors~u~and~v`then~u.v=\left| u \right|\left| v \right|\cos \theta \]

OpenStudy (anonymous):

I have no idea what that means...

OpenStudy (anonymous):

I can't get the ||u|| part.. What do those things mean?

OpenStudy (anonymous):

u=8i+7j v=9i+7j \[u.v=\left( 8i+7j \right).\left( 9i+7j \right)=\left( 8 \right)\left( 9 \right)+\left( 7 \right)\left( 7\right)=72+49=121\] \[\left| u \right|=\sqrt{8^2+7^2}=\sqrt{64+49}=\sqrt{113}=3\sqrt{13}\] \[\left| v \right|=\sqrt{9^2+7^2}=\sqrt{81+49}=\sqrt{130}\] \[121=3\sqrt{13}\sqrt{130}\cos \theta=3(13)\sqrt{10}\cos \theta ,\] \[\cos \theta=\frac{ 121 }{39\sqrt{10} }=\frac{ 121\sqrt{10} }{ 390 }=?\]

OpenStudy (anonymous):

.9811 Then do I do cos^-1(.9811)?

OpenStudy (anonymous):

Nope.. That's not it.

OpenStudy (anonymous):

@surjithayer

OpenStudy (whpalmer4):

@surjithayer \[\sqrt{113}\ne3\sqrt{13}\]because \(9*13= 117, \text{ not } 113\)

OpenStudy (anonymous):

Oh. -_-

OpenStudy (anonymous):

YES YOU ARE CORRECT \[\cos \theta=\frac{ 121 }{\sqrt{113}\sqrt{130} }=0.9983,\theta=\cos^{-1} 0.9983 =3.3 ~degree\]

OpenStudy (whpalmer4):

That'll derail any of the computations downstream from there, of course...

OpenStudy (anonymous):

sorry for that.

OpenStudy (anonymous):

Okay, got it! You're fine, hun. We're only human. :3 Thank you! God bless you!

OpenStudy (whpalmer4):

oh, good, you're available, I'm in the middle of cooking dinner, and was dreading fixing the problem :-)

OpenStudy (whpalmer4):

by the way, if you want to make a degree symbol, use ^\circ

OpenStudy (anonymous):

in the equation

OpenStudy (anonymous):

or on the computer.

OpenStudy (anonymous):

Are you talking to me or him?

OpenStudy (whpalmer4):

I meant if you wanted to have something like \[3.3^\circ\] \ [3.3^\circ\ ] (without the spaces) will do the trick

OpenStudy (anonymous):

\[\left[ 3.5^\circ \right] \]

OpenStudy (anonymous):

Oh, okay. :3 Thanks!